The mass of a beaker is 5.833 g. After 3.00 mL of a certain pure liquid, is added into the beaker, the combined mass of the beaker and the liquid sample is 8.613 g. From the data, what is the measured density of the liquid?


1
Expert's answer
2022-02-02T06:48:02-0500

Mass of liquid"=8.613-5.833"

"=2.78g"


Volume of liquid"=3ml"

"=3cm^3"


Density "=\frac{mass}{volume}=\frac{2.78}{3}=0.9267g/cm^3"